Let h(x) = ln(x^2 + 1) be a function defined on (−∞,∞). Find the equation of the tangent line to the curve of h(x) at x = 1. Use the exact values only
Answer #1

Given that hix) = ln (x²+1) (-oo, oo) + at x=1 we know that The equation of the tangent 1 at point (oxo, Yol line of the funcand h(x) = en (x²+1) ln (2) hell In(it) In(60²+1) C hill In (2) (x, y) - = (+, 1n121) by en y- ln(2) (1). (x - 1) x-1 y- ln(2